OpenStudy (anonymous):

sigma n=1 to infinity (10n +3)/(n2^n) converge or diverge ?

OpenStudy (anonymous):

converges like crazy

OpenStudy (anonymous):

thank you could you please tell me the first step to solve it @satellite73

OpenStudy (anonymous):

oh sorry i would use the eyeball test \[2^n\] is exponential grows way faster than the polynomial in the numerator but you could use the ratio test probably if you want something more math teachery

OpenStudy (anonymous):

\[a_n=\frac{10n+3}{n2^n}\] \[a_{n+1}=\frac{10(n+1)+3}{(n+1)2^{n+1}}\]

OpenStudy (anonymous):

find \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}\]